OpenStudy (anonymous):

how does the integral of [1/(x-1)(x+3)]dx equal [2/3(ln((x-1)/(4-x))+ ln2]?

zepdrix (zepdrix):

Hi there :) welcome to OpenStudy! \[\Large\rm \int\limits \frac{1}{(x-1)(x+3)}dx\]This requires Partial Fraction Decomposition.

zepdrix (zepdrix):

\[\Large\rm \frac{1}{(x-1)(x+3)}=\frac{A}{(x-1)}+\frac{B}{(x+3)}\]Multiplying through by the denominator on the left gives us,\[\Large\rm 1=A(x+3)+B(x-1)\]

zepdrix (zepdrix):

Any confusion up to that point? :o

OpenStudy (anonymous):

no, I understand so far,:)

OpenStudy (anonymous):

What else do I do?

OpenStudy (anonymous):

Do you know how to solve the partial fraction?

OpenStudy (anonymous):

yes I know how to solve the partial fraction and when I solve it I get something like 1-4(ln (x-1)/(x+3))

OpenStudy (anonymous):

well, for the partial fractions I get (1/(4(x-1))-(1/(4(x+3)))

OpenStudy (anonymous):

Were your values for A and B in the partial fractions 1/4 and -1/4 respectively?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok that's correct. Now if we integrate that we get 1/4ln(x-1)-1/4ln(x+3) So your integral is correct i believe...

OpenStudy (anonymous):

Wolfram also gives that. https://www.wolframalpha.com/input/?i=integrate+1%2F%28%28x-1%29%28x%2B3%29%29

OpenStudy (anonymous):

yes, but the answer is not available. the correct answer is (2/3)(ln((x-1)/(4-x))+ln2)

OpenStudy (anonymous):

the problem says something like: we find A,B so that 2/(x-1)(4-x)=(A/(x-1))+(B/(4-x))

OpenStudy (anonymous):

please somebody help me find the answer?:(

OpenStudy (anonymous):

that really confuses me, I don't know how they got that

zepdrix (zepdrix):

That looks completely different from this problem 0_o

zepdrix (zepdrix):

You sure you're matching up the correct answer? D:

OpenStudy (anonymous):

yes, they were the corrections posted after I answered the problem and I didnt understand how they came up with it, I can try to upload a picture?

zepdrix (zepdrix):

That would probably be a good idea :3

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

sorry about the bad quality, I was not able to screenshot it so I took a picture with my phone

zepdrix (zepdrix):

So here is the answer key 0_o Hmmmm... no picture of the problem? :c

OpenStudy (anonymous):

oh okay yes Ill do that

OpenStudy (anonymous):

How come the problem is different to the working steps shown in the photo?

OpenStudy (accessdenied):

\( \displaystyle \int \dfrac{1}{(x-1)\color{#00aa00}{(x+3)}} \ dx \ \color{red}\ne \ \dfrac{2}{3} \ln \left( \dfrac{x-1}{\color{#aa0000}{4-x}} \right) + c \) @ilymartin , I think the issue here is a typographical error. The 'correct' answer provided does not match the integral you began with; it appears to stem from a different integrand. Since you say this was provided after answering the question and should match, I suspect the issue was in the service and not in your Math. :)